### v1

# APB: Accelerating Distributed Long-Context Inference by Passing Compressed Context Blocks across GPUs

Abstract
While long-context inference is crucial for advancing large language model (LLM) applications, its prefill speed remains a significant bottleneck. Current approaches, including sequence parallelism strategies and compute reduction through approximate attention mechanisms, still fall short of delivering optimal inference efficiency. This hinders scaling the inputs to longer sequences and processing long-context queries in a timely manner. To address this, we introduce APB, an efficient long-context inference framework that leverages multi-host approximate attention to enhance prefill speed by reducing compute and enhancing parallelism simultaneously. APB introduces a communication mechanism for essential key-value pairs within a sequence parallelism framework, enabling a faster inference speed while maintaining task performance. We implement APB by incorporating a tailored FlashAttn kernel alongside optimized distribution strategies, supporting diverse models and parallelism configurations. APB achieves speedups of up to 9.2, 4.2, and 1.6 compared with FlashAttn, RingAttn, and StarAttn, respectively, without any observable task performance degradation. We provide the implementation and experiment code of APB in https://github.com/thunlp/APB.
APB: Accelerating Distributed Long-Context Inference by
Passing Compressed Context Blocks across GPUs
Yuxiang Huang1††thanks: indicates equal contribution., Mingye Li2∗††thanks: Work done during internship at TsinghuaNLP., Xu Han1††thanks: indicates corresponding authors., Chaojun Xiao1, Weilin Zhao1, Ao Sun3, Hao Zhou4, Jie Zhou4, Zhiyuan Liu1, Maosong Sun1 1NLP Group, DCST, IAI, BNRIST, Tsinghua University, Beijing, China. 2Department of CS&T, Central South University, Changsha, China. 3BUPT, Beijing, China. 4Pattern Recognition Center, WeChat Al, Tencent Inc. huang-yx21@mails.tsinghua.edu.cn, lmy2004@csu.edu.cn, han-xu@tsinghua.edu.cn, xiaocj20@mails.tsinghua.edu.cn
1 Introduction
Large language models (LLMs) (OpenAI, 2024; Anthropic, 2024; DeepSeek-AI, 2024) have demonstrated unprecedented proficiencies, pushing the boundaries of artificial intelligence research and practical applications. Recent advancements are not only transforming usage paradigms but also empowering intelligent systems such as LLM-based agents (Li, 2025; Qin et al., 2024; Zhao et al., 2024), robotics (Zeng et al., 2023; Kim et al., 2024), and prompting methodologies (Chu et al., 2023; Sahoo et al., 2024). These systems often rely on extended context inference. To address the growing demand for longer inputs, contemporary foundation models have been increasingly designed to support extended context lengths. For instance, Llama-3.1 (Dubey et al., 2024) supports up to 128K tokens, Claude-3.5 (Anthropic, 2024) extends input capacity to 200K tokens, and MiniMax-01 (Li et al., 2025) can even process input sequences up to 4M tokens.
As context lengths grow, the quadratic computational cost of attention makes single-GPU inference both infeasible and inefficient for LLMs. To address this, various optimizations aim to enhance parallelism or reduce compute. Sequence parallelism (Li et al., 2023), which aims to enhance parallelism, partitions the sequence across devices (termed as hosts) and significantly improves the prefill speed, especially for extremely long inputs (Figure 1). However, the overall computations remain unchanged to ensure the accuracy of the attention results. On the other hand, approximate attention mechanisms (Zhang et al., 2024e; Li et al., 2024b; Jiang et al., 2024), which compute only those elements selected from the attention matrix, accelerate inference by reducing compute but face scalability challenges and performance degradation when processing longer inputs. To this end, designing an approximate attention mechanism that fits in sequence parallelism frameworks offers a promising way to further enhance efficiency, particularly in accelerating long-context prefill. However, designing such systems demands system and algorithm optimizations to address the key challenges.
Challenge 1: Localized Attention Pruning. Existing widely-used approximate attention mechanisms, such as H2O (Zhang et al., 2024e) and SnapKV (Li et al., 2024b), typically depend on full sequence information, such as the attention scores computed over the entire sequence, to prune the redundant compute of the attention scores. This requirement directly conflicts with the distributed architecture of sequence parallelism, where individual hosts only maintain partial context visibility without heavy host-to-host communication.
Challenge 2: Multi-host Scalability. Traditional sequence parallelism approaches face inherent scalability limitations due to model-architectural constraints and performance degradation risks. While sequence parallelism with attention head splitting for computation (Jacobs et al., 2023) offers substantial throughput improvements, its scalability remains fundamentally bounded by the fixed number of attention heads. Existing solutions that simply combine approximate attention with sequence parallelism, such as StarAttn (Acharya et al., 2024), suffer from progressive performance degradation when the number of hosts increases, as these solutions merely extend Xiao et al. (2024b) to multiple hosts with a large proportion of invisible context.
To address these challenges, we propose APB, a distributed inference framework designed to leverage approximate attention to reduce redundant compute and communication overhead. For localized attention pruning, APB introduces a local key-value (KV) cache compression technique that operates independently on each host, eliminating the need for a global sequence view to prune redundant attention compute. For multi-host scalability, APB ensures that each host processes all attention heads within its local context and selectively shares compressed critical context across hosts. This design enables APB to maintain stable model performance even as the number of hosts scales up.
We implement APB using a customized Flash-Attn Dao (2024) kernel and an optimized distribution framework, enabling efficient scaling across diverse sequence lengths and multiple hosts. Comprehensive evaluations demonstrate that APB achieves an excellent trade-off between inference speed and model performance across a variety of tasks. Additionally, APB is compatible with various model sizes and distribution settings, making it a robust solution for scalable distributed inference. APB achieves speedups of up to 9.2, 4.2, and 1.6 compared with FlashAttn, RingAttn Li et al. (2023), and StarAttn, respectively, without any observable performance degradation.
2 Related Works
Existing approaches to the efficient long-context inference of LLMs focus on system and algorithm optimizations. Please refer to the surveys Zhao et al. (2023); Wan et al. (2024) for more about LLMs.
System Optimizations. The efficiency of long-context inference can be enhanced by leveraging hardware architectures and preserving accurate LLM computations. Hardware-aware algorithms, such as FlashAttn (Dao et al., 2022; Dao, 2024; Shah et al., 2024), utilize matrix tiling to optimize GPU memory usage and boost inference speed. Additionally, efficient parallelism methods support longer sequences with a higher processing speed. RingAttn (Li et al., 2023) splits long sequences across multiple hosts using ring-style communication to preserve accurate attention computation, while Ulysses (Jacobs et al., 2023) distributes attention heads across hosts to reduce communication overhead. Other parallelism strategies (Narayanan et al., 2021; Huang et al., 2019; Rajbhandari et al., 2020; Sun et al., 2025; DeepSeek-AI, 2024) enable longer sequences on larger models, and the mixture of various distribution strategies (Ren et al., 2021; Sun et al., 2024a) can further enhance long-context efficiency. However, most parallelism approaches are tailored for training and lack optimization for inference. Offloading-based methods Xiao et al. (2024a); Sun et al. (2024b); Lee et al. (2024); Chen et al. (2025) leverage hierarchical memory systems to reduce hardware requirements by storing redundant KV cache in CPU memory and recalling only a small part to GPU memory.
Algorithm Optimizations. The burden of long-context inference can also be mitigated through algorithm optimizations. KV cache-centric optimization reduces the size of the KV cache, enabling faster decoding speed and lower GPU memory usage. Cache eviction methods Zhang et al. (2024e); Li et al. (2024b); Yao et al. (2024); Huang et al. (2024) discard irrelevant or unimportant KV cache, alleviating the memory bottlenecks. Quantization techniques Liu et al. (2024); Hooper et al. (2024); Zhang et al. (2024a); He et al. (2024) decrease the memory footprint by utilizing low-bit representations to store the KV cache. Merging methods Cai et al. (2024); Wang et al. (2024); Zhang et al. (2024c, d) consolidate redundant KV cache units across sequences or layers to alleviate storage overheads. Sparse mechanisms Zaheer et al. (2020); Beltagy et al. (2020); Lou et al. (2024) decrease inference latency by reducing the attention computational load. Specifically, approaches like MInference Jiang et al. (2024) and FastGen Ge et al. (2024) assign specific patterns to attention heads, accelerating inference by computing only elements selected from the attention matrix. Moreover, algorithm optimizations can complement system optimizations. StarAttn Acharya et al. (2024) and APE Yang et al. (2025) linearize attention complexity by dividing context into parallelized blocks, but they struggle with tasks that require inter-context dependencies. More details on these algorithm optimizations are elaborated in Li et al. (2024a) and Shi et al. (2024). Apart from KV cache optimizations on Transformer-based Vaswani (2017) LLMs, approaches altering backbone architectures can also enhance long-context inference efficiency. Typically, RNN-based models (Gu and Dao, 2024; Dao and Gu, 2024) and hybrid models (Lieber et al., 2025; Dong et al., 2024; Li et al., 2025) reduce the computation complexity of long-context inference.
3 Methodology
3.1 Preliminaries
We primarily introduce the Transformer architecture and the prefill-decoding pattern.
Transformers. Transformer-based LLMs take a sequence as input. We denote the hidden state of the token at the -th layer as , where represents the embeddings of the input sequence. Assuming the model has layers, each layer consists of an attention block and a feedforward network (FFN) block. For each attention head in an attention block, the attention score at layer is calculated as follows, where represents the attention mask.
where , , and . The hidden size of each head is denoted by . Then, the attention scores are passed through the FFN: .
Prefill and Decoding. The procedure of an LLM processing an input can be divided into two stages: prefill and decoding. During the prefill stage, the model computes the KV cache for the input sequence and predicts the first answer token . This process can be formalized as . After the prefill stage, the model decodes one token at a time using the KV cache of previous tokens, which can be written as . The complexities of the prefill stage and a single step of decoding are and , respectively.
3.2 Framework
We adopt sequence parallelism with approximate attention to accelerate the prefill of long-context inputs. The framework of APB is illustrated in Figure 2. The inference of APB can be split into 4 stages: context splitting, block compression, communication, and computation.
Context Splitting. Given a long-context input , we first split the input sequence into a document and a query , i.e., and . Since , we primarily focus on optimizing the prefill speed of the document.We apply accurate attention with online softmax for processing the query and decoding the answer. The document is evenly split across all hosts, with an anchor block prepended at the start of each local context block on every host to preserve its visibility to the document’s initial part.
Block Compression. To reduce the attention compute and communication overhead, we first perform KV cache compression using a compressor on each host to shorten the block length and only retain the essential KV cache units for inter-host communication in the next step. These compressed blocks are crucial for capturing long-distance semantic dependency and maintaining task performance by reducing attention errors.
Communication. We design a specialized communication pattern to gather the compressed context blocks on each host for the awareness of previous essential context. After the communication, we form the passing blocks to encapsulate the essential KV cache units passed from the previous hosts.
Computation. Finally, each host obtains a context layout containing the local context block, an anchor block, and a passing block. We perform attention using a modified attention mask, executed with a specialized FlashAttn (Dao, 2024) kernel.
Overall, APB is designed to communicate and compute only the most critical context by aggressively compressing the KV cache within a sequence parallelism framework. It introduces an algorithm-aware, distributed system-level optimization specifically tailored to its unique communication pattern and approximate attention mechanism.
3.3 Context Splitting
At the beginning of APB inference, we first split the document across hosts, with each host holding a partial sequence with length , denoted as blocks . Each block is prepended with an anchor block at the front. Inspired by StarAttn Acharya et al. (2024), we adopt the anchor block containing the first tokens of the input sequence. We also embed the query in the front of the anchor block, to provide more query-related information for the compressors to identify the essential cache units. Therefore, the anchor block can be formally written as . The positions assigned to the tokens in anchor blocks are the starting positions, i.e. . Notably, we apply much smaller anchor blocks compared to StarAttn, where or in APB but in StarAttn. Formally, each host except the first one holds a sequence of . The first host only holds without anchor block.
3.4 Block Compression
Before starting the attention computation on the -th host, we compress the KV cache of into using the compressors . As calculating full attention scores or full KV cache on a single host is infeasible, existing methods like H2O Zhang et al. (2024e) and SnapKV Li et al. (2024b) are incompatible in this scenario. The need to identify important KV cache units without a global view aligns with the design of Locret Huang et al. (2024), where importance scores are based on each token’s query and KV cache. We implement the compressors as Locret’s retaining heads , which are small MLPs trained on additional long-context SFT data. These MLPs take as input and output importance scores for the corresponding cache units, reflecting the influence towards the calculation of subsequent KV cache. Formally, the compressed context block on host with passing length is defined as where and .
3.5 Communication
To acquire all the compressed context blocks sent from each host after block compression, we apply an AllGather communication on the compressed KV cache across all the hosts by . Subsequently, the passing block is constructed by concatenating all compressed context blocks from previous hosts, i.e. . We simply ignore the compressed context blocks sent by subsequent hosts.
3.6 Computation
After the communication and construction of the passing blocks, each host acquires the context layout required for attention computation. We place the passing block between the anchor block and the local context block on each host to perform the attention as Equation 2. The passing blocks are discarded after the attention calculation and do not participate in the FFN calculation. is the modified attention mask. We implement the attention calculation in a modified FlashAttn Dao (2024) kernel with only the attention mask changed.
We present the complete inference procedure of APB in Algorithm 1. Running approximate attention of the document is detailed in Algorithm 2, while the accurate attention with online softmax is presented in Algorithm 3.
| Method | R.PassKey | R.Number | R.KV | E.Sum | E.QA | E.MC | E.Dia | Z.QA | C.Debug | M.Find | Avg. |
| Llama-3.1-8B-instruct | |||||||||||
| FullAttn | 100.00 | 99.49 | 51.00 | 30.59 | 29.04 | 63.76 | 11.00 | 36.18 | 24.62 | 28.82 | 47.45 |
| MInference | 98.47 | 98.81 | 17.40 | 30.06 | 26.42 | 55.46 | 12.50 | 36.51 | 28.17 | 32.29 | 43.61 |
| StarAttn | 100.00 | 98.98 | 40.60 | 30.55 | 30.66 | 61.57 | 15.00 | 36.26 | 26.65 | 24.57 | 46.48 |
| APB | 100.00 | 98.81 | 81.8 | 30.63 | 29.25 | 63.32 | 11.00 | 36.99 | 30.46 | 26.86 | 50.91 |
| Qwen-2.5-14B-instruct | |||||||||||
| FullAttn | 100.00 | 100.00 | 17.80 | 27.80 | 10.40 | 52.84 | 28.00 | 10.21 | 38.07 | 42.57 | 42.68 |
| MInference | 100.00 | 100.00 | 5.20 | 25.63 | 12.04 | 61.57 | 16.50 | 11.06 | 41.62 | 48.57 | 42.22 |
| StarAttn | 100.00 | 100.00 | 18.00 | 27.48 | 12.44 | 62.88 | 23.50 | 11.03 | 43.40 | 37.71 | 43.69 |
| APB | 100.00 | 100.00 | 62.20 | 26.51 | 12.74 | 66.38 | 25.00 | 11.22 | 37.31 | 35.71 | 47.34 |
| Yi-34B-200K | |||||||||||
| FullAttn | 100.00 | 100.00 | 49.00 | 5.83 | 17.57 | 47.60 | 2.00 | 18.77 | 25.13 | 28.00 | 39.39 |
| MInference | 100.00 | 100.00 | 52.40 | 7.19 | 20.07 | 63.32 | 2.50 | 25.40 | 28.17 | 31.40 | 43.05 |
| StarAttn | 100.00 | 100.00 | 24.60 | 4.38 | 19.71 | 60.26 | 0.00 | 23.40 | 28.17 | 21.14 | 38.17 |
| APB | 100.00 | 100.00 | 65.20 | 5.96 | 19.81 | 62.45 | 1.00 | 25.42 | 27.92 | 30.00 | 43.78 |
| Method | SG1 | SG2 | SG3 | MK1 | MK2 | MK3 | MV | MQ | VT | CWE | FWE | QA1 | QA2 | Avg. |
| Llama-3.1-8B-instruct | ||||||||||||||
| FullAttn | 99.40 | 99.80 | 99.60 | 98.20 | 87.60 | 67.00 | 94.65 | 98.00 | 60.98 | 71.40 | 72.20 | 78.20 | 41.6 | 82.20 |
| MInference | 100.00 | 98.60 | 99.00 | 95.40 | 58.20 | 23.80 | 84.35 | 95.70 | 66.40 | 45.94 | 74.67 | 67.80 | 38.80 | 72.97 |
| StarAttn | 100.00 | 99.60 | 99.60 | 95.80 | 73.60 | 53.00 | 72.80 | 94.45 | 59.40 | 65.72 | 76.53 | 67.40 | 41.00 | 76.84 |
| APB | 100.00 | 100.00 | 99.80 | 85.60 | 91.00 | 89.00 | 95.05 | 96.40 | 51.96 | 63.82 | 77.33 | 70.00 | 41.20 | 81.63 |
| Qwen-2.5-14B-instruct | ||||||||||||||
| FullAttn | 100.00 | 99.20 | 99.80 | 94.20 | 47.80 | 27.20 | 75.10 | 94.60 | 89.52 | 93.88 | 76.13 | 63.20 | 43.40 | 77.23 |
| MInference | 100.00 | 99.60 | 100.00 | 89.60 | 32.60 | 6.94 | 61.25 | 93.29 | 89.44 | 84.36 | 77.33 | 56.60 | 44.40 | 71.95 |
| StarAttn | 100.00 | 99.20 | 97.40 | 80.00 | 38.00 | 20.20 | 48.75 | 77.10 | 82.44 | 93.86 | 77.53 | 52.80 | 38.00 | 69.64 |
| APB | 100.00 | 100.00 | 99.80 | 93.00 | 87.20 | 85.80 | 66.55 | 93.00 | 96.40 | 94.94 | 76.33 | 60.00 | 55.60 | 85.28 |
| Yi-34B-200K | ||||||||||||||
| FullAttn | 100.00 | 100.00 | 99.60 | 95.20 | 76.00 | 55.40 | 92.10 | 97.05 | 85.56 | 51.84 | 84.27 | 65.20 | 50.00 | 80.94 |
| MInference | 100.00 | 100.00 | 100.00 | 96.00 | 87.00 | 59.60 | 90.15 | 96.05 | 73.68 | 79.20 | 84.87 | 67.40 | 49.40 | 83.33 |
| StarAttn | 100.00 | 100.00 | 99.60 | 93.20 | 78.20 | 48.00 | 77.05 | 88.35 | 83.96 | 80.58 | 78.33 | 61.20 | 48.80 | 79.79 |
| APB | 100.00 | 100.00 | 100.00 | 92.80 | 88.00 | 92.60 | 86.05 | 96.45 | 60.48 | 88.46 | 84.53 | 63.80 | 52.20 | 85.03 |
4 Experiments
In this section, we present our benchmark results to evaluate APB, aiming at addressing the following research questions: (Q1) Can APB obtain higher end-to-end prefill speed without significant performance degradation? (Q2) How effective and efficient is APB on various context length? (Q3) Is the design of each component in APB effective?
4.1 Experimental Setup
In this section, we briefly introduce experimental settings, and more details are in Appendix B.
Benchmarks. We evaluate APB against selected baselines on two long-context evaluation benchmarks: Bench (Zhang et al., 2024b) and RULER Hsieh et al. (2024). Bench is a benchmark with an average context length exceeding 100K tokens, encompassing a mix of both synthetic and real-world tasks. Following MInference Jiang et al. (2024), we utilize 10 tasks, excluding Math.Calc and Code.Run due to their difficulty, which leads to failures for all methods. RULER is a benchmark featuring a variety of synthetic tasks with a controllable context length. We evaluate all 13 tasks of the benchmark. We provide the mapping of task abbreviations to the original task name in Appendix D. Additional experiments and results are placed in Appendix E.
Models. For the end-to-end benchmark (Q1), we test APB with the baselines on the instruct versions of Llama-3.1-8B, Qwen-2.5-14B, and Yi-34B-200K. When benchmarking across various input lengths (Q2), we utilize Llama-3-8B-1M111https://huggingface.co/gradientai/Llama-3-8B-Instruct-Gradient-1048k, which supports up to 1M input tokens.
Metrics. When evaluating performance on the benchmark, we directly adopt the original performance metrics for each task. For speed measurement, we define inference speed as the total number of tokens in both the prefill and decoding stages, divided by the total inference time, which includes both the prefill and decoding stages.
Baselines. Existing methods for accelerating long-context inference can be categorized into four types based on their use of sequence parallelism and approximate attention mechanisms. We select FlashAttn, Ulysses, RingAttn, MInference, and StarAttn as our baselines. FlashAttn does not use sequence parallelism and maintains accurate attention computation. Ulysses and RingAttn are two major approaches for performing accurate attention with sequence parallelism, each with a different communication design. MInference is an approximate attention mechanism that accelerates long-context inference by applying sparse attention without sequence parallelism. StarAttn combines sequence parallelism and approximate attention by prepending an anchor block to each context block and removing all communication. Further details on the baselines can be found in Appendix C.
4.2 End-to-End Results
To address Q1, we conduct an end-to-end benchmark focusing on the overall performance and long-context processing speed by comparing APB with all the baselines on both Bench and RULER.
APB demonstrates superior task performance in both real-world scenarios (Bench in Table 1) and synthetic benchmarks (RULER in Table 2), compared to all the selected baselines. In certain tasks, APB even surpasses FullAttn. In contrast, MInference and StarAttn exhibit noticeable performance degradation relative to FullAttn. Notably, APB significantly enhances performance in complex context retrieval tasks, such as R.KV in Bench and MK3 in RULER. This improvement can be attributed to APB’s ability to trim noisy context within each block by our compressor, thereby constructing cleaner passing blocks with reduced noise and enabling more accurate context retrieval.
In terms of inference speed, APB also demonstrates significant advantages over existing methods. As illustrated in Figure 3, methods without sequence parallelism, such as FlashAttn and MInference, exhibit slower inference speeds. In contrast, sequence parallelism-based methods, including RingAttn and Ulysses, achieve speedups ranging from 3 to 10 compared to FlashAttn. By incorporating approximate attention, StarAttn achieves a faster inference speed than RingAttn, though its improvement over Ulysses remains limited. APB, on the other hand, delivers at least a 25% speed improvement over StarAttn without any performance decay.
Overall, APB establishes the best speed-performance trade-off, outperforming all competing methods in both efficiency and effectiveness.
4.3 Benchmarking Across Different Lengths
We evaluate APB and baselines with input lengths ranging from 32K to 512K to address Q2. All methods are tested on the RULER benchmark, where we report task performance, inference speed, and the compute under each input length in Figure 4.
As shown in Figure 44 (a), APB consistently outperforms all the methods across all input lengths in terms of performance. Additionally, APB achieves the fastest inference speed among all methods, as depicted in Figure 44 (b). The speed advantage of APB remains humble for shorter input sequences. However, as the input length increases, APB’s advantage becomes more pronounced, demonstrating a substantial speedup over all competitors at a 512K input length. Notably, both StarAttn and APB show an increase in inference speed from 32K to 128K inputs, whereas other methods continue to slow down. Since StarAttn and APB both incorporate approximate attention mechanisms, the overall compute is reduced, delaying the point where compute becomes the dominant bottleneck instead of memory access. Figure 44 (c) further explains APB’s superior speed, as it yields significantly less compute than FullAttn and StarAttn, especially for longer inputs. Further discussion on the compute of each method can be referred to Table 6.
4.4 Ablation Studies
Regarding Q3, we conduct ablation studies and focus on the following key points.
| No. | A | P | Q | E.MC | |
| 0 | ✓ | ✓ | ✓ | 72.00 | |
| 1 | ✓ | ✓ | ✗ | 68.00 | |
| 2 | ✓ | ✓ | Rd. | ✓ | 66.00 |
| 3 | ✓ | ✓ | Rd. | ✗ | 66.00 |
| 4 | ✓ | ✗ | Rd. | ✓ | 64.00 |
| 5 | ✓ | ✗ | Rd. | ✗ | 62.00 |
| 6 | ✗ | ✓ | ✗ | 28.00 | |
| 7 | ✗ | ✓ | Rd. | ✗ | 28.00 |
| 8 | ✗ | ✗ | Rd. | ✗ | 22.00 |
Each Component’s Contribution. We focus on four major components in the APB framework: the anchor block, the passing block, selecting important KV pairs via retaining heads, and embedding the query at the beginning of the anchor block. We conduct an ablation study on these four components. For the anchor block and passing block, they can be removed in the corresponding settings. To ablate the implementation of the compressors , we compare retaining heads with random selectors picking the same number of KV pairs. When the embedded query is removed, the anchor block contains only the beginning of the document.
The ablation results in Table 3 show that all four components are effective, as removing any of them leads to performance degradation (No. 1, 2, 4, 6). The anchor block is the most critical component, as its removal causes task failure (No. 6, 7). The passing block is also essential for good performance, as a performance drop of more than 8% occurs in No. 4 and 5. The retaining heads are crucial for selecting important KV pairs to pass to subsequent hosts, as randomly selecting KV pairs results in poorer performance (No. 2). Embedding the query in the anchor block must be used in conjunction with the retaining heads, as activating these components separately does not improve performance (No. 1, 2, 3). Above all, when used together, they yield significantly better performance (No. 0).
Wall-time Breakdown Analysis. We begin by analyzing the prefill and decoding time, as detailed in Appendix E.1. Our findings indicate that prefill dominates the overall processing time, while decoding time is negligible in long-context tasks.
Next, we decompose the prefill time into seven components: the projection of QKV, the calculation of retaining heads, the communication time, the attention computation, the projection of O, the FFN computation, and others. The wall-time of each component is shown in Figure 5. Methods without sequence parallelism (FlashAttn and MInference) exhibit substantial inference time, whereas sequence parallelism proves effective in reducing the time of both attention and FFN computations.
Compared to existing sequence parallelism methods, APB achieves even lower attention time without incurring significant overhead for retaining heads and communication. When compared with RingAttn and Ulysses, although the introduction of the anchor blocks increases the compute cost for FFN, the additional time is smaller than the reduction in attention time, thus resulting in a faster inference speed. APB outperforming StarAttn is due to smaller anchor blocks, whose overhead in the FFN time is significantly lower.
Distributed Settings. We evaluate the effectiveness of APB across various distributed settings. As mentioned in StarAttn (Acharya et al., 2024), when the context is split across more hosts, significant performance degradation occurs due to the increasing invisibility of the middle context. However, APB addresses this issue by using passing blocks, allowing each host to share the most essential KV pairs with subsequent hosts.
We test APB and StarAttn across hosts, and the results are shown in Table 4. When the sequence is long (128K tokens), both APB and StarAttn exhibit improved performance as the sequence parallelism size increases. With a shorter input length (32K tokens), the block size on each host is smaller, and APB demonstrates consistently stronger and more stable performance across all settings, while StarAttn shows significant performance loss as the host count increases.
| Length | Method | ||||
| 128K | APB | 84.00 | 92.00 | 92.00 | 92.00 |
| StarAttn | 82.00 | 84.00 | 86.00 | 88.00 | |
| 32K | APB | 92.00 | 92.00 | 94.00 | 94.00 |
| StarAttn | 94.00 | 92.00 | 92.00 | 84.00 |
5 Conclusion
We propose APB, a distributed long-context inference framework based on approximate attention, achieving speedups up to 9.2, 4.2, and 1.6 compared to FlashAttn, RingAttn, and StarAttn, respectively, without observable performance degradation. APB constructs the passing blocks with minimal communication cost by sending only core KV pairs among hosts. The experimental results on Bench and RULER, using various models and sequence lengths, demonstrate that APB delivers faster inference speeds while maintaining or exceeding performance. Furthermore, APB is adaptable to diverse distribution configurations and models of varying sizes. Next, we aim to accelerate the decoding process in APB, where the KV cache is distributed across different hosts.
Limitations
As APB is specifically optimized to minimize the prefill time for extremely long inputs, it is less effective for processing shorter inputs, particularly those under 32K tokens. When applying APB to short inputs, the optimal distributed setting is to run the inference on a single host, as the computation can be effectively parallelized within a single GPU. In such cases, APB falls back to vanilla FlashAttn.
Appendix A Pseudocode of APB
Here, we present the complete inference process of APB in Algorithm 1.
For initialization, the input sequence is first split into the document and query (line 1). Each host then takes a partial context based on the host index (line 1) and sets the anchor block (lines 1–1). To begin prefill, the embeddings of the anchor block A and the context block are computed (line 1). During the prefill stage (lines 1–1), we employ Algorithm 2 to perform sequence parallelism-aware approximate attention. In the decoding stage (lines 1–1), we directly apply accurate attention with online softmax (Algorithm 3), as introduced in StarAttn (Acharya et al., 2024) as stage-2, to generate new tokens. Finally, the generated tokens are returned (line 1).
The prefill stage function of APB is described in Algorithm 2. Before performing the attention calculation, the KV cache of the current host is compressed using the retaining heads (line 2–2). Next, two AllGather communications are performed on the compressed K and V caches, allowing each host to obtain a full view of the compressed KV cache from all hosts (line 2–2). Subsequently, the passing block is formed by concatenating the compressed KV caches sent from the previous hosts (line 2), followed by the modified attention mechanism (line 2). After the attention calculation, the passing blocks are discarded, and only the attention scores corresponding to the anchor block and the context block held by the current host are passed through the FFN (line 2).
The decoding function of APB, also known as stage-2 of StarAttn, is described in Algorithm 3. In the decoding stage, all hosts hold the hidden states H of the new token. The Q, K, and V of the new token are computed simultaneously on each host (line 3). For all hosts except the last one, the attention score is calculated by the new token’s query with the KV cache held by the current host (line 3). For the last host, the local KV cache is first concatenated with the new KV, and then attention calculation is performed with the query (line 3). A Gather communication is performed to collect and lse calculated by each host (line 3), which are then merged to obtain the global attention score A (line 3). Notably, the A on each host should be identical. Finally, A is passed through the FFN simultaneously on each host, generating identical on all hosts (line 3).
Appendix B Reproducibility Settings
B.1 Training of the Compressor
We use the retraining heads proposed in Huang et al. (2024) as our compressor. Following Huang et al. (2024), we set the intermediate size of the retaining heads to 1024. We utilize the first 3000 samples from the LongAlign (Bai et al., 2024) dataset to generate the training labels, and we train the retaining heads with the a frozen model backbone for 3000 steps in 1 epoch. The batch size set is 1 and the maximum input length is set to 10240. We use AdamW as the optimizer and the learning rate is set to 5e-4, with set to 0.9 and set to 0.95. We apply a linear scheduler whose number of warmup steps is set to 300. Notably, the training loss of the retaining head consists of a regression loss and a smoothing loss, following the setting of Huang et al. (2024), and the balance factor is set to 0.0025. The gradient clipping value is set to 0.5 to avoid gradient explosion.
| 32K | 64K | 128K | 256K | 512K | |
| 4K | 8K | 16K | 32K | 64K | |
| 1K | 2K | 4K | 8K | 8K | |
| 0.5K | 1K | 2K | 4K | 8K |
B.2 Inference Hyperparameters
Here, we elaborate on the details of the inference hyperparameters used in Section 4.2, 4.3, and 4.4.
B.2.1 End-to-End Benchmark (Section 4.2)
In the end-to-end benchmark, we first test the performance of 3 baselines with the performance of our method. Since methods without approximate attention do not alter the computational outcome of the attention mechanism, we take the result of Ulysses as the performance benchmark for FullAttn. Then, we evaluate the inference speed of all baselines along with APB .
Performance Evaluation. In the Bench experiments, we evaluate the performance of each method by running all the data from each task. Details of the tasks can be found in Zhang et al. (2024b). When conducting the RULER experiments, we generate 500 test samples for each RULER task to evaluate the performance, with details introduced in Hsieh et al. (2024). We report the performance of each method under a 128K input length for both benchmarks. For methods with sequence parallelism, we set the sequence parallel size to 8. We use a single machine with 8 GPUs for Llama-3.1-8B and Qwen-2.5-14B, while for Yi-34B, due to its large model size, we employ two machines with layers evenly distributed between them. All the answers are generated with greedy decoding, i.e. with a temperature set to 0. For StarAttn, we set both the block length and the anchor length to 16K. For MInference, we use the official head configuration for Llama-3.1-8B and generate the configurations for Qwen-2.5-14B and Yi-34B based on the first data entry of SG1 from RULER. For APB, we set anchor length to 4K and passing length to 2K. We place the query tokens after the system prompt to embed the query within the anchor blocks without disrupting the model’s chat template.
Speed Evaluation. We run the first 20 samples for each task and average the inference speed. To reflect the speed of LLMs in processing a query, we define the inference speed as follows.
We keep all the hyperparameters and distribution settings same as the performance evaluation.
B.2.2 Benchmarking Across Input Lengths (Section 4.3)
For Section 4.3, we evaluate the performance, speed, and report the compute of all the methods across different input lengths using the Llama-3.1-8B model and the RULER benchmark. We run 50 samples per task for performance evaluation, while for speed testing, we run 20 samples per task. For StarAttn, we set both the block size and the anchor length to , where is the input length and is the sequence parallelism size. For APB, we use the hyperparameters listed in Table 5. We align other hyperparameters to the end-to-end benchmark. We visualize the compute of each method following Table 6.
B.2.3 Ablation Studies (Section 4.4)
Here, we elaborate on the settings in Section 4.4.
Each Component’s Contribution. We conduct the ablation study of each component using Llama-3.1-8B. We conduct experiments of different settings on E.MC from Bench using the first 50 samples to assess the impact of different combinations on performance. We use the hyperparameters of .
Wall-time Breakdown Analysis. The prefill and decoding time is tested with Llama-3.1-8B on SG1 from RULER. For the breakdown analysis of the prefill time, we run all the methods on synthetic random input. All the experiments are conducted under a 128K input length, and the hyperparameters are kept the same as in previous experiments.
Distributed Settings. We test the performance under different host numbers (sequence parallelism size) with Llama-3.1-8B. We set across and evaluate the performance of StarAttn and APB on E.MC from Bench. We set the sequence length to 32K and 128K, and test on the first 50 samples.
B.3 System Environment
All training and evaluation are conducted on workstations equipped with 8 NVIDIA A800-80GB GPUs and 104 Intel® Xeon® Platinum 8470 CPUs, running CentOS Linux 7 (Core). The GPUs within each machine are interconnected using third-generation NVLink technology, while cross-machine communication is facilitated via HDR InfiniBand. We perform the training of the retaining head and all experiments—except for FlashAttn and MInference tests—on 8 GPUs. The experiments for FlashAttn and MInference are conducted on a single GPU.
Appendix C Baselines
In this section, we briefly introduce the design principles of each baseline.
FlashAttn. Dao (2024) introduces an attention computation method with hardware awareness that leverages the architecture of modern GPUs. FlashAttn applies tiling techniques to the calculation of attention scores and fully utilizes the high-bandwidth memory to accelerate computation. FlashAttn is an accurate attention method that preserves the computation results, primarily focusing on accelerating attention on a single GPU.
MInference. As introduced in Jiang et al. (2024), three approximate attention patterns can be utilized to approximate FullAttn computation. MInference first searches for the head configuration by assigning an approximate pattern to each head. During inference, only a limited number of attention score entries are calculated, accelerating the prefill by reducing the computation in attention. MInference focuses on applying approximate attention without sequence parallelism.
RingAttn. Introduced by Li et al. (2023), RingAttn distributes the context across multiple GPUs (hosts). Online softmax is applied to the attention calculation, where each host computes the partial attention score passed by the previous host, using local context for rounds ( is the number of hosts). The core idea is to overlap the communication of passing partial attention scores with the attention calculation. RingAttn is primarily used for training LLMs with extremely long contexts but is not optimized for inference.
Ulysses. Similar to RingAttn, Ulysses (Jacobs et al., 2023) was also introduced for long-context training. Ulysses applies sequence parallelism, with each host holding a partial context. During attention calculation, three AlltoAll communications on Q, K, and V are performed to distribute the full context of specific heads to the corresponding host. After the calculation, another AlltoAll communication is conducted on the attention score to revert the distribution from splitting heads to splitting the sequence.
StarAttn. Introduced by Acharya et al. (2024), StarAttn is a pioneer in combining approximate attention with sequence parallelism. StarAttn introduces anchor blocks, which contain the initial tokens of the input sequence and are the same size as the blocks on each host. Each host calculates only the partial attention between the anchor block and the current context block, with no communication required. Although StarAttn reduces communication and synchronization, its large anchor blocks introduce overhead in the FFN, limiting the speedup gains.
| Method | FLOPs / forward |
| FullAttn | |
| StarAttn | |
| APB |
Here, we present the computation per forward call for each method in Table 6. Notably, since FlashAttn, RingAttn, and Ulysses are FullAttn methods, they share the same computation formula. We do not include MInference, as its computation depends on the search results of the head configurations. We provide a visualization of Table 6 in Figure 44 (c).
Appendix D Task Abbreviations
Here, we provide a mapping between the benchmark task abbreviations and their full names used in the main text in Table 7 and 8.
| Abbr. | Full Name |
| R.PassKey | Retrieve.PassKey |
| R.Number | Retrieve.Number |
| R.KV | Retrieve.KV |
| E.Sum | En.Sum |
| E.QA | En.QA |
| E.Dia | En.Dia |
| Z.QA | Zh.QA |
| C.Debug | Code.Debug |
| M.Find | Math.Find |
| Abbr. | Full Name |
| SG1 | Single NIAH 1 |
| SG2 | Single NIAH 2 |
| SG3 | Single NIAH 3 |
| MK1 | Multi-keys NIAH 1 |
| MK2 | Multi-keys NIAH 2 |
| MK3 | Multi-keys NIAH 3 |
| MV | Multi-values NIAH |
| MQ | Multi-queries NIAH |
| VT | Variable Tracking |
| CWE | Common Words Extraction |
| FWE | Frequent Words Extraction |
| QA1 | Question Answering 1 |
| QA2 | Question Answering 2 |
Appendix E Further Ablation Studies and Detailed Experimental Results
Here, we represent the further ablation studies in Section E.1 and E.2. We list all experimental data that are not included in the main text in Section E.3.
E.1 Breakdown Analysis: Prefill and Decoding
We conduct a wall-time breakdown analysis to further investigate how APB achieves a higher inference speed. First, we measure the prefill and decoding times for each method, and the results shown in Figure 6 reveal that the prefill stage is the bottleneck, while the decoding stage takes up significantly less time. Since APB optimizes the prefill stage, it is able to address the most time-consuming part of the process.
E.2 Hyperparameter Stability
There are two hyperparameters introduced in the APB framework: the length of the anchor block , and the passing length . In this study, we examine the sensitivity of these two parameters by measuring the performance on the E.QA task from Bench for various lengths of and .
In this experiment, we use Llama-3.1-8B as the model backbone. We test on the first 50 samples and set the sequence length to 128K. We present the results with .
| Method | R.PassKey | R.Number | R.KV | E.Sum | E.QA | E.MC | E.Dia | Z.QA | C.Debug | M.Find | Avg. |
| Llama-3.1-8B-instruct | |||||||||||
| FlashAttn | 4056 | 3998 | 3530 | 609 | 3985 | 3924 | 4280 | 4137 | 3959 | 5209 | 3769 |
| Ulysses | 30999 | 30426 | 25266 | 3423 | 27616 | 25860 | 29654 | 25332 | 31289 | 39967 | 26983 |
| RingAttn | 20440 | 20218 | 17507 | 3082 | 19476 | 18577 | 20826 | 18779 | 20092 | 25956 | 18495 |
| MInference | 7910 | 7686 | 5259 | 832 | 6186 | 7238 | 7016 | 6399 | 7719 | 8027 | 6427 |
| StarAttn | 33968 | 32632 | 25840 | 2605 | 27896 | 31986 | 32267 | 27266 | 31942 | 37051 | 28345 |
| APB | 44995 | 43111 | 32086 | 2848 | 32930 | 41040 | 36553 | 32057 | 44780 | 45117 | 35552 |
| Qwen-2.5-14B-instruct | |||||||||||
| FlashAttn | 2155 | 2082 | 1736 | 706 | 1929 | 1758 | 1863 | 1882 | 2131 | 2347 | 1859 |
| Ulysses | 17214 | 16461 | 12623 | 1574 | 13374 | 13398 | 14092 | 11131 | 16932 | 18460 | 13526 |
| RingAttn | 11056 | 10589 | 8658 | 1314 | 10140 | 9539 | 9982 | 8977 | 10919 | 11676 | 9285 |
| MInference | 3527 | 3305 | 2244 | 618 | 2645 | 2498 | 2480 | 2222 | 3669 | 3785 | 2699 |
| StarAttn | 18793 | 17892 | 13695 | 1641 | 14604 | 13874 | 14462 | 12453 | 18640 | 19640 | 14569 |
| APB | 25218 | 23643 | 17192 | 1592 | 16875 | 17937 | 17267 | 13303 | 25134 | 25929 | 18409 |
| Yi-34B-200K | |||||||||||
| FlashAttn | 1041 | 1089 | 907 | 372 | 1110 | 1137 | 1099 | 1125 | 1111 | 1208 | 1030 |
| Ulysses | 9192 | 8820 | 7383 | 1369 | 8523 | 8388 | 8841 | 8374 | 9016 | 9901 | 7981 |
| RingAttn | 6203 | 6032 | 5296 | 1281 | 6021 | 5785 | 5951 | 6024 | 5989 | 6552 | 5513 |
| MInference | 1732 | 1637 | 1199 | 400 | 1327 | 1760 | 1523 | 1062 | 1766 | 1806 | 1421 |
| StarAttn | 9039 | 8675 | 7284 | 1893 | 8419 | 9007 | 8866 | 7877 | 9046 | 9373 | 7948 |
| APB | 12441 | 9404 | 2059 | 10798 | 12726 | 12233 | 9327 | 12744 | 9328 | 12744 | 10624 |
The experimental results shown in Figure 7 indicate that both and are stable. While there is a slight performance improvement with larger , the variation remains insignificant. The performance change with varying is also minimal. This suggests that it is not necessary to tune and delicately, making APB easy to use.
E.3 The Detailed Experimantal Results
Here, we present the detailed experimental results to extensively illustrate each figure in the main text, Section E.1 and E.2. In Section 4.2, the detailed experimental results of Figure 3 are provided in Table 9 and Table 12. Table 9 provides the accurate speed for all methods on Bench, while the speed is the average value across all tasks. By contrast, Table 12 reports the precise speed measurements for all methods on RULER, with the values similarly averaged over all tasks. The performance of each method in Figure 3 is provided in Table 1 and Table 2. In Section 4.3, Figure 14 and Figure 15 present the detailed results corresponding to Figure 4. Specifically, Table 14 aligns with Figure 44 (a) while Table 15 corresponds to Figure 44 (b). In Section 4.4, the detailed experimental results of Figure 5 are provided in Table 13. In Section E.1, we add the breakdown analysis of the prefill and decoding wall-time for various methods, while the results are illustrated in Figure 6. The detailed experimental results are provided in Table 10. In Section 1, the detailed experimental results of Figure 1 are provided in Table 11.
| Prefill Time | Decoding Time | |
| FlashAttn | 30137.03 | 422.31 |
| Ulysses | 4028.66 | 263.52 |
| RingAttn | 6317.03 | 185.33 |
| MInference | 18067.32 | 2275.61 |
| StarAttn | 3556.60 | 208.23 |
| APB | 2554.77 | 284.21 |
| 32K | 64K | 128K | 256K | 512K | 1024K | |
| FlashAttn | 3.46 | 9.51 | 30.01 | OOM | OOM | OOM |
| Ulysses | 0.50 | 1.30 | 3.95 | 13.49 | 49.55 | OOM |
| RingAttn | 0.72 | 2.00 | 6.34 | 21.80 | 81.62 | OOM |
| MInference | 4.95 | 8.72 | 15.16 | OOM | OOM | OOM |
| StarAttn | 0.67 | 1.43 | 3.50 | 9.63 | 30.43 | OOM |
| APB | 0.33 | 0.50 | 0.94 | 2.64 | 6.48 | 19.40 |
| Method | SG1 | SG2 | SG3 | MK1 | MK2 | MK3 | MV | MQ | VT | CWE | FWE | QA1 | QA2 | Avg. |
| Llama-3.1-8B-instruct | ||||||||||||||
| FlashAttn | 4277 | 4282 | 4062 | 4317 | 4293 | 4145 | 3908 | 3899 | 3847 | 4203 | 4056 | 4361 | 4382 | 4156 |
| Ulysses | 30138 | 29537 | 25214 | 29611 | 29542 | 24749 | 20721 | 20212 | 27411 | 19994 | 21721 | 29547 | 31609 | 26154 |
| RingAttn | 19642 | 19521 | 17318 | 19760 | 19622 | 17298 | 14582 | 15072 | 18716 | 14731 | 16244 | 20022 | 19862 | 17876 |
| MInference | 7660 | 7579 | 7555 | 7362 | 5780 | 5201 | 4509 | 3245 | 7720 | 7148 | 6544 | 5604 | 3617 | 6117 |
| StarAttn | 31636 | 32532 | 25900 | 32436 | 32314 | 25378 | 17494 | 20770 | 29319 | 22738 | 22901 | 19696 | 33661 | 26675 |
| APB | 43930 | 44504 | 33106 | 44502 | 43717 | 33882 | 31328 | 25349 | 39680 | 25580 | 27828 | 42931 | 45664 | 37077 |
| Qwen-2.5-14B-instruct | ||||||||||||||
| FlashAttn | 2175 | 1958 | 1819 | 2054 | 1925 | 1438 | 1426 | 1315 | 1882 | 1352 | 1982 | 2038 | 2045 | 1801 |
| Ulysses | 16221 | 13968 | 13761 | 15047 | 14713 | 11345 | 10552 | 9863 | 14312 | 11420 | 15174 | 15642 | 15464 | 13652 |
| RingAttn | 10571 | 9558 | 9322 | 10070 | 9892 | 7771 | 7791 | 7487 | 9816 | 8267 | 10505 | 10464 | 10358 | 9375 |
| MInference | 3391 | 2097 | 2637 | 2408 | 2389 | 1476 | 1845 | 1634 | 2721 | 1465 | 2756 | 3134 | 3010 | 2382 |
| StarAttn | 18226 | 16314 | 15403 | 18285 | 15941 | 11292 | 13469 | 9292 | 15276 | 11039 | 16059 | 16118 | 15965 | 14821 |
| APB | 24576 | 21588 | 19580 | 20732 | 23305 | 14292 | 15638 | 10786 | 18601 | 13469 | 19301 | 21758 | 20339 | 18767 |
| Yi-34B-200K | ||||||||||||||
| FlashAttn | 1194 | 1146 | 1122 | 1147 | 1182 | 1144 | 1078 | 1099 | 1174 | 1140 | 1224 | 1167 | 1186 | 1154 |
| Ulysses | 8466 | 5513 | 4427 | 5110 | 8659 | 7498 | 6144 | 6674 | 8152 | 7083 | 8099 | 8130 | 8130 | 7083 |
| RingAttn | 4478 | 3895 | 3915 | 5158 | 5814 | 5286 | 4657 | 4862 | 5615 | 5092 | 5685 | 5620 | 5625 | 5054 |
| MInference | 1644 | 1279 | 1199 | 1340 | 1585 | 1293 | 1012 | 1101 | 1450 | 1210 | 1361 | 1484 | 1461 | 1340 |
| StarAttn | 8554 | 7569 | 6939 | 7690 | 8713 | 7442 | 6933 | 6704 | 8108 | 7086 | 7826 | 8160 | 8151 | 7683 |
| APB | 11989 | 10176 | 8902 | 10030 | 12061 | 9768 | 7910 | 8396 | 10252 | 9135 | 10264 | 10990 | 10777 | 10050 |
| QKV Projection | Retaining Head | Communication | Attention | O Projection | FFN | Others | Transformer Block | |
| FlashAttn | 25.33 | – | – | 664.01 | 17.42 | 201.44 | 32.67 | 940.86 |
| Ulysses | 3.31 | – | 3.90 | 84.53 | 2.27 | 25.88 | 4.62 | 124.51 |
| RingAttn | 3.21 | – | 18.40 | 152.12 | 2.09 | 24.40 | 4.62 | 205.19 |
| MInference | 24.45 | – | – | 281.39 | 15.88 | 201.56 | 40.80 | 564.07 |
| StarAttn | 6.67 | – | – | 41.84 | 4.29 | 50.01 | 8.56 | 111.37 |
| APB | 4.01 | 1.72 | 0.62 | 34.07 | 2.67 | 30.76 | 6.33 | 80.18 |
| Method | SG1 | SG2 | SG3 | MK1 | MK2 | MK3 | MV | MQ | VT | CWE | FWE | QA1 | QA2 | Avg. |
| K | ||||||||||||||
| FullAttn | 100.00 | 100.00 | 98.00 | 100.00 | 96.00 | 82.00 | 97.00 | 98.50 | 92.00 | 40.20 | 88.00 | 82.00 | 64.00 | 87.52 |
| MInference | 100.00 | 100.00 | 100.00 | 100.00 | 96.00 | 76.00 | 95.50 | 99.00 | 90.40 | 59.40 | 88.00 | 80.00 | 62.00 | 88.18 |
| StarAttn | 100.00 | 100.00 | 100.00 | 96.00 | 98.00 | 96.00 | 83.50 | 93.50 | 88.80 | 76.20 | 90.67 | 78.00 | 62.00 | 89.44 |
| APB | 100.00 | 98.00 | 100.00 | 100.00 | 98.00 | 100.00 | 89.50 | 98.50 | 86.40 | 78.00 | 90.00 | 74.00 | 60.00 | 90.18 |
| K | ||||||||||||||
| FullAttn | 100.00 | 100.00 | 98.00 | 100.00 | 98.00 | 56.00 | 99.00 | 98.00 | 84.40 | 1.20 | 78.67 | 68.00 | 54.00 | 79.64 |
| MInference | 100.00 | 100.00 | 100.00 | 100.00 | 98.00 | 54.00 | 97.50 | 99.50 | 78.00 | 6.20 | 81.33 | 64.00 | 58.00 | 79.73 |
| StarAttn | 100.00 | 94.00 | 100.00 | 96.00 | 96.00 | 86.00 | 81.50 | 94.50 | 82.40 | 16.00 | 82.00 | 70.00 | 52.00 | 80.80 |
| APB | 100.00 | 96.00 | 100.00 | 98.00 | 98.00 | 100.00 | 95.00 | 97.50 | 80.40 | 11.40 | 85.33 | 72.00 | 58.00 | 83.97 |
| K | ||||||||||||||
| FullAttn | 100.00 | 100.00 | 100.00 | 98.00 | 100.00 | 36.00 | 98.50 | 95.50 | 77.20 | 0.00 | 72.00 | 68.00 | 46.00 | 76.25 |
| MInference | 100.00 | 100.00 | 100.00 | 100.00 | 100.00 | 46.00 | 96.50 | 99.00 | 74.80 | 0.20 | 80.00 | 76.00 | 54.00 | 78.96 |
| StarAttn | 100.00 | 100.00 | 100.00 | 96.00 | 96.00 | 90.00 | 79.50 | 90.50 | 84.40 | 0.20 | 72.00 | 68.00 | 52.00 | 79.12 |
| APB | 100.00 | 100.00 | 100.00 | 98.00 | 94.00 | 98.00 | 97.00 | 98.00 | 73.20 | 0.20 | 70.67 | 68.00 | 52.00 | 80.70 |
| K | ||||||||||||||
| FullAttn | 100.00 | 100.00 | 96.00 | 94.00 | 97.22 | 22.00 | 92.50 | 95.00 | 64.00 | 0.60 | 76.67 | 78.00 | 44.00 | 73.85 |
| MInference | 100.00 | 100.00 | 100.00 | 96.00 | 94.00 | 46.00 | 91.85 | 95.50 | 77.20 | 0.20 | 79.33 | 72.00 | 48.00 | 76.93 |
| StarAttn | 100.00 | 96.00 | 100.00 | 94.00 | 86.00 | 66.00 | 78.50 | 83.00 | 70.00 | 0.20 | 80.67 | 68.00 | 42.00 | 74.18 |
| APB | 100.00 | 96.00 | 100.00 | 94.00 | 94.00 | 94.00 | 93.00 | 98.00 | 47.60 | 0.40 | 78.00 | 70.00 | 46.00 | 77.77 |
| K | ||||||||||||||
| FullAttn | 98.00 | 98.00 | 100.00 | 94.00 | 76.00 | 10.00 | 90.50 | 96.00 | 46.80 | 0.60 | 86.67 | 70.00 | 46.00 | 70.20 |
| MInference | 100.00 | 100.00 | 100.00 | 98.00 | 76.00 | 12.00 | 85.50 | 91.00 | 77.20 | 0.40 | 83.33 | 62.00 | 42.00 | 71.34 |
| StarAttn | 100.00 | 98.00 | 100.00 | 88.00 | 74.00 | 20.00 | 71.50 | 80.00 | 45.20 | 0.40 | 88.00 | 60.00 | 38.00 | 66.39 |
| APB | 100.00 | 100.00 | 100.00 | 90.00 | 88.80 | 80.00 | 88.80 | 95.00 | 25.20 | 0.80 | 83.33 | 72.00 | 44.00 | 74.33 |
| Method | SG1 | SG2 | SG3 | MK1 | MK2 | MK3 | MV | MQ | VT | CWE | FWE | QA1 | QA2 | Avg. |
| K | ||||||||||||||
| FlashAttn | 8955 | 8764 | 8754 | 7369 | 8818 | 7534 | 6259 | 6751 | 8005 | 6614 | 8200 | 8333 | 8121 | 7883 |
| Ulysses | 42479 | 40140 | 22599 | 39735 | 45166 | 23239 | 14300 | 15644 | 27447 | 15465 | 26981 | 32685 | 25030 | 28532 |
| RingAttn | 31432 | 32855 | 16703 | 29762 | 32016 | 19901 | 14227 | 15243 | 21899 | 12687 | 20532 | 26173 | 22230 | 22743 |
| MInference | 4940 | 4714 | 3116 | 4644 | 4725 | 3018 | 1771 | 2097 | 3765 | 2094 | 3608 | 3874 | 3657 | 3540 |
| StarAttn | 33909 | 34256 | 17630 | 34951 | 36137 | 20604 | 14362 | 12704 | 23505 | 14008 | 23792 | 27637 | 24730 | 24479 |
| APB | 50778 | 43160 | 18264 | 41896 | 45641 | 21149 | 24893 | 13909 | 30407 | 13457 | 25016 | 29156 | 30863 | 29892 |
| K | ||||||||||||||
| FlashAttn | 6647 | 6664 | 6026 | 6606 | 6627 | 6208 | 5006 | 5526 | 6334 | 5619 | 6411 | 6344 | 6468 | 6191 |
| Ulysses | 38672 | 37904 | 22892 | 36890 | 38257 | 24925 | 13190 | 18511 | 28011 | 17790 | 27585 | 28760 | 30551 | 27995 |
| RingAttn | 27484 | 27209 | 20289 | 27089 | 27621 | 21167 | 11823 | 16726 | 22120 | 15897 | 22340 | 23076 | 23959 | 22062 |
| MInference | 5531 | 5437 | 3954 | 5527 | 5156 | 4012 | 2069 | 2531 | 4609 | 3130 | 4364 | 4539 | 4868 | 4287 |
| StarAttn | 35871 | 35840 | 21658 | 33590 | 33620 | 23213 | 17342 | 16487 | 28760 | 19524 | 27363 | 26143 | 26949 | 26643 |
| APB | 48822 | 46944 | 32005 | 50754 | 51244 | 29646 | 19554 | 19655 | 30925 | 21003 | 28469 | 29516 | 30056 | 33738 |
| K | ||||||||||||||
| FlashAttn | 4262 | 4233 | 3984 | 4285 | 4209 | 4062 | 3463 | 3766 | 4161 | 3910 | 4330 | 4184 | 4266 | 4086 |
| Ulysses | 30195 | 29300 | 25446 | 30020 | 30070 | 25720 | 16690 | 20717 | 28030 | 22484 | 28064 | 26920 | 26951 | 26200 |
| RingAttn | 20013 | 19221 | 17226 | 19559 | 19565 | 17264 | 13162 | 15344 | 18501 | 16042 | 18622 | 18715 | 18447 | 17822 |
| MInference | 5389 | 5385 | 4459 | 5427 | 5097 | 4597 | 2275 | 3167 | 5149 | 3762 | 4851 | 4789 | 4734 | 4545 |
| StarAttn | 34734 | 34416 | 29086 | 34171 | 33984 | 28468 | 19095 | 20370 | 31682 | 25456 | 31313 | 31339 | 30678 | 29600 |
| APB | 46644 | 46110 | 36421 | 42653 | 46052 | 35107 | 18480 | 27517 | 40831 | 30995 | 39476 | 40254 | 37931 | 37575 |
| K | ||||||||||||||
| FlashAttn | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM |
| Ulysses | 18365 | 18295 | 16987 | 18385 | 18265 | 17231 | 12866 | 15530 | 17689 | 16117 | 17894 | 17710 | 17735 | 17159 |
| RingAttn | 11847 | 11748 | 11335 | 11777 | 11849 | 11431 | 9325 | 10654 | 11642 | 10925 | 11799 | 11661 | 11663 | 11358 |
| MInference | 5046 | 5223 | 4262 | 5027 | 4744 | 4411 | 2315 | 2711 | 4930 | 3705 | 4682 | 4519 | 4559 | 4318 |
| StarAttn | 26455 | 26174 | 23985 | 26096 | 26304 | 24073 | 18374 | 18399 | 25383 | 22303 | 24674 | 25413 | 24953 | 24045 |
| APB | 34126 | 33613 | 29983 | 34000 | 33914 | 30404 | 19710 | 25390 | 32728 | 26596 | 31437 | 31574 | 32469 | 30457 |
| K | ||||||||||||||
| FlashAttn | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM |
| Ulysses | 10327 | 10249 | 10089 | 10299 | 10317 | 9993 | 8910 | 9334 | 10324 | 9797 | 10458 | 10192 | 10259 | 10042 |
| RingAttn | 6395 | 6337 | 6246 | 6331 | 6383 | 6253 | 5783 | 5937 | 6342 | 6138 | 6508 | 6332 | 6332 | 6255 |
| MInference | 4796 | 4720 | 3999 | 4468 | 4465 | 4134 | 2261 | 2314 | 4585 | 3613 | 4562 | 4367 | 4332 | 4047 |
| StarAttn | 16841 | 16661 | 16130 | 16556 | 16746 | 15988 | 13936 | 14365 | 16969 | 15737 | 16969 | 16798 | 16600 | 16177 |
| APB | 28100 | 27614 | 25938 | 28315 | 28306 | 25773 | 19435 | 22188 | 27541 | 24470 | 27484 | 27087 | 27082 | 26102 |